Your in a closed circuit race; 1/5 of the bikers are in front of you and 5/6 of the bikers are in back of you; what the is number of bikers are in the race?
Answer by khwang(438)

If totally, there are bikers,by the given condition, we have
x/5 + 1 + 5 x/6 = x.
Or (6x + 25x)/30 - x = -1,
Or x/30 = -1
So, x = -30

However, the total numbers of bikers cannot be negative.
Hence,there is no solution.
[Note,since 1/5 + 5/6 = 31/30 >1 ,while 1 means the all bikers.
So, front bikers + back bikers > all bikers is impossible.]